A cubical lantern needs wire as frame. If each edge is 25cm, what is the least length of wire needed if joints are soldered?

Answer by Theo(13342)

you are talking about a cube with sides that have a length of 25 cm.
the volume of the cube is 25^3 = 15625 cubic cm.
the surface area of the cube is 6 * 25^2 = 3750 square cm.
there are 12 edges to the cube with each edge requiring 25 cm of wire which makes a total of 12 * 25 = 300 cm of wire.
a picture of what i think your cube lantern will look like and the number of edges that require wire is shown below.
$$$$